Operation Summary for April 14, 2010 - April 16, 2010 |
Wednesday April 14 The day shift began with Operations delivering beam to MiniBooNE, Main Injector was off due to a problem with a 3 way cooling valve and the Tevatron (TeV) cooldown at F2 was progressing at a slower rate than expected. During the day shift, the 3 way cooling valve was placed in manual and beam returned to Pbar and NuMI at 9:12 AM. At 1:54 PM, cryogenic experts reported that the plug at F2 had been cleared and that cooldown should be completed in 6 to 8 hours. Experts began changing the MIRF 3 way valve at 2:10 PM, no beam was delivered to Pbar and NuMI the rest of the shift. During the evening shift, both D0 and CDF made accesses lasting about 90 minutes. The MIRF 3 way valve was replaced, but technicians were unsuccessful in calibrating it. The valve was placed in manual and beam returned to Pbar and NuMI at 5:30 PM. The Tev was cold at 8:30 PM and experts began startup. The startup was delayed at 11:21 PM when the Tev abort link went down at F0, operators unsuccessfully tried to reset the abort, and an expert was called in to investigate. Thursday April 15 The midnight shift began with Operations monitoring the stash (301.5E10), with Pbar stacking (44.9E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, controls experts found a faulty reset switch on the interlock chassis at F0 and Tevatron checkout resumed at 2:12 AM. Operations put the TeV into HEP shot setup. Operations established store 7743 at 5:02 AM with an average initial luminosity (AIL) of 280.77E30. Operators noticed that the Q849 MI-8 line shunt was showing erratic behavior thus causing beam excursions in the beamline, an expert was contacted. During the day shift, Store 7743 collided all shift. MI beam was off for 40 minutes while a power supply tech changed the regulator for Q849 MI-8 line shunt. LRF2 tripped off and a bad power supply was replaced, which resulted in 60 minutes of downtime. During the evening shift, Store 7743 collided all shift. Operations performed Linac and Booster tuning to reduce losses.Friday April 16 The midnight shift began with Operations monitoring store 7743 (21.5E30) and the stash (432.6E10), with Pbar stacking (22.9E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7743 at 12:30 AM. D0 made a 60 minute access. Opertions put the TeV into HEP shot setup and established store 7746 at 4:00 AM with an initial luminosity of 378.16E30 (a new record). Operations noticed that the Linac buncher appears to have an intermittent short, experts will address the problem on the day shiftThe Plans for Friday and the Weekend The plan for today is to repair the Linac buncher short and calibrate the MIRF 3 way valve, then stack, stash, and store through the weekend. |
|
More Information Comments and Suggestions |
| last modified 04/16/2010 email Fermilab |
| Security, Privacy, Legal | |